

********************************************************************
UARS WEEKLY STATUS REPORT - 20 Dec 2002
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE was powered only during daylight for two days to
support the yaw maneuver. Otherwise, solar data were collected on all
available orbital tracking periods.
Happy Winter SOLSTICE from the SOLSTICE team!
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** Happy Holidays from PEM *****
The PEM web site has moved. It has a new URL. Access the PEM web site
using http://wwwpem.space.swri.edu. The PEM data from the GUVI conjunction
periods has moved to http://wwwpem.space.swri.edu/guvi-pem.html. It was
discovered that this web site was not updated properly with the latest PEM
data plots. SwRI has fixed this problem.
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 1994039 has been
reprocessed. HEPSA data reprocessing continues but has been slowed
temporarily due to the Holidays. PEM HEPSA data from 1994040 through
1994159 is waiting for DAAC archive.
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes when both instruments saw aurora.
PEM has seen no aurora during the second conjunction period. Due to a
spacecraft anomaly, no PEM data was collected during the third conjunction
period.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of Educational Outreach and
the Science Traceability Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 December 2002 (Orbit 61465, GMT Day 341) through
13 December 2002 (Orbit 61571, GMT Day 347).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 345/2145 61539 POWER OFF AXIS-1
PEM 347/2257 61570 POWER OFF
The beta angle ranged between the angles of 32.3 degrees to 16.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 18 December 2002 (DOY 352).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C. Additional
battery performance data for each battery on 06 December and 13
December 2002 are provided in Appendix D.
The clock error ranged from -10.0 to -3.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 01
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Three (3) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began testing the OBC image from the OBC crash with the
simulator this week. Problems with the UTTS simulator caused a one
day setback in testing. The converted load files were all accepted by
the simulator. The FOT then successfully started the OBC. Before
starting the OBC the FOT did a precursory check of all the residual
OBC values. Nothing out of the ordinary was seen.
The OBC was then initialized. All processors were turned on and
initialized with no problems. The attitude began to settle in with
the Earth Sensor enabled. The FOT plans on doing some more detailed
testing, but initial signs indicate nothing out of the ordinary that
might have caused the OBC to crash.
MEETINGS
The FOT attended a TDRS Naming Convention Meeting to discuss how the
new TDRS naming convention would be handled with our current CMS
version.
The FOT also attended a SOI Planning meeting to discuss possible
changes to the curriculum for the upcoming spring semester.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 December 2002 (DOY 352).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 December 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
61539 345/2145 02-485 PEM POWER OFF AXIS-1
61569 347/2114 02-487 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
61570 347/2257 02-488 PEM POWER OFF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 December 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 December 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
61569 347/0128 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 December 2002
06 December 2002 - GMT Day 340 Beta = 33.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -156.8/-179.2 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-33.6 +5.79/+4.51 +27.2/-14.4 25.1 1.040 21.2
3 +22.4/-44.8 +5.15/+3.87 +26.4/-13.6 25.1 1.030 20.7
13 December 2002 - GMT Day 347 Beta = 16.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -156.8/-179.2 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +44.8/-22.4 +6.43/+5.14 +29.6/-14.8 24.2 1.030 22.0
3 +22.4/-50.4 +5.14/+3.87 +28.8/-15.2 24.2 1.020 21.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 December 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-061 FOT 61512 344/0124 31907 33 Second Data Dropout
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-058 FOT 61113 317/1341 N/A NEG ACQ / UFLTR ERROR
02-060 FOT 61421 338/0057 N/A PSU Cmd Clock Timeout
02-061 FOT 61512 344/0124 31907 33 Second Data Dropout